OpenStudy (fernandez555):

***medal+award*** help please! Given the functions f(x) = -2x - 1 and g(x) = -3x + 4, which operation results in the smallest coefficient on the x term? A.Two operations result in the same coefficient B. Subtraction C.Multiplication D.Addition

OpenStudy (perl):

so i assume you are looking at f - g , f + g , f * g

OpenStudy (fernandez555):

@perl yeap

OpenStudy (perl):

f*g = ( -2x -1 ) ( -3x + 4) = 6x^2 +3x -8x - 4 = 6x^2 -5x - 4 f+ g = (-2x - 1 ) + (-3x + 4) = -5x - 3 . so the answer is two of them

OpenStudy (perl):

f*g = ( -2x -1 ) ( -3x + 4) = 6x^2 +3x -8x - 4 = 6x^2 -5x - 4 f+ g = (-2x - 1 ) + (-3x + 4) = -5x - 3 f - g = ( -2 x - 1 ) - (-3x + 4) = -2x - 1 + 3x - 4 = x - 5 so we have two expression with -5x

OpenStudy (fernandez555):

@perl thank you!